Determine the equilibrium constant for the reaction NH4 + (aq) + OH– (aq) ⇌ H2O(ℓ) + NH3 (aq) given that the Ka of NH4+ is 5.68×10–10 .
Answer #1

Ka of NH4+ is:

NH4+ (aq) <—> NH3(aq) + H+(aq)

So,

Ka = [NH3][H+]/[NH4+]

= 5.68*10^-10

The dissociation of water:

H2O(l) <—> H+(aq) + OH-(aq)

Kw =[H+][OH-]/H2O

= 1.0*10^-14

The given equation is:

NH4+(aq) + OH-(aq) <—> H2O(l) + NH3(aq)

Kc = [H2O][NH3]/[NH4+][OH-]

= Ka/Kw

= (5.68*10^-10)/ (1.0*10^-14)

= 5.68*10^4

Answer: 5.68*10^4